Shift summary for Owl, Sept. 9, 1999

Angela Bellavance (bellavan@rice.edu)
Thu, 9 Sep 1999 07:04:29 -0500 (CDT)

Shift Summary for OWL shift, 09/Sep/99

On Shift: Rick K., Angela B

Hours of beam delivered at all intensities: 6.1/7.0
Hours of beam delivered at nominal intensity: 6.1
Hours of E832 data (special runs not included): 4.1
Hours of special runs (laser scans, muon runs...): 1.6

Accounting of lost time:
putting regenerator back in, changing tapes

========================================================================
Runs ended this shift:
Run 14472 - 3.25M ECAL tags - BEAM_832_KE3HOG2 (100 MeV ptkick)
Run 14473 - 410 K ECAL tags - BEAM_832_NOTHRESH (100 MeV ptkick)
Run 14474 - ongoing - BEAM_832 (400 MeV ptkick)

Tapes brought to FCC this shift: QKB143, QKE037-QKE045
(including tapes for online split)

========================================================================
Accesses during shift:
23 min - Elizabeth and Angela to put regenerator back in position.

========================================================================
Overview of shift:
special runs (ke3hog2 and no csi threshold) and regular datataking

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
14472 11 00:05 Running BEAM_832_KE3HOG2 with reg OUT when beam
returns. Elizabeth paged to come in for putting
reg back in.

Noticed that 1 KHz spill clock not working.
This was traced to bad 622 coinc unit at the
end of this run; cables moved to adjacent section.

21 Ktevana errors - KFRENDs.
22 00:20 Pipeline errors, stop run after 3.3M ECAL tags.
This makes 4M ECAL tags when combined with earlier
run.

00:26 Elizabeth and Angela go on access to put reg. in.
REGMOVE process started and reg HV restored.

14473 01 00:31 Start BEAM_832_NOTHRESH run with no CsI threshold
for B02-Ke3. Time in spill working again.

00:49 Return from access.
23 01:02 Beam returns.

To run smoothly, set L1 pre-scales
so that L1-B07[V*mu] rate = 5.5 KHz and L1-B02
rate = 0.8 KHz. Ran for almost one hour with NO
pipeline problems. If the dummy B07 rate is
reduced, we get lots of pipeline errors.

28 01:09 Beam gone for one spill.
01:45 Reached our goal of 400K ECAL tags; run stopped.
01:55 SY energy conservation in effect - 1 spill

14474 1 02:02 Begin regular data-taking run after switching
tapes and turning the analysis magnet back on
(NEG).
5 02:07 AZLAT crate 53 goes FULL
02:30 Minor abort in TeV (no noticable beam loss).
145 05:15 Beam goes away. Ch.13 claims SY energy
conservation.
05:24 Ch.13 msg: Investigating beam loss.
183 06:07 Beam returns!
189 06:15 SY energy cons. for one spill.
06:49 MCR calls, says beam going away in 10 minutes.
07:00 Hand off to day shift.

========================================================================
Follow up Notes on Previous Shift

Finished special runs, except KE3HOG2 with reg in at 4E12
[wait for more discussion ... do we really need it ?]

========================================================================
Items to Follow up on Next Shift

Laser scan when beam goes away.
Day shift needs to contact DaveN to have him replace the center island
ArEth bottles.

For Rick, Angela

_______
Angela Bellavance work: (630) 840-4144
bellavan@rice.edu home: (630) 548-3412
http://www.owlnet.rice.edu/~bellavan/ page: (630) 722-6861
"It has yet to be proven that intelligence has any survival value."
- Arthur C. Clarke